An order is written for phenytoin IV. A loading dose of 15mg/kg is to be infused at 0.5mg/kg/min for a 33 pound child. The Pharmacy has phenytoin injection solution 50mg/ml in a 5ml vial in stock. The pharmacist will put the dose into 50 ml NS. Over how many minutes should the dose be administered? Round to the nearest whole number.

For Pharmacists (Disclaimer-for calculation practices only) 15. An order is written for phenytoin IV. A loading dose of 15mg/THIRD TIME POSTING, PLEASE HELP ME!!
Answer #1

Child weight is 33 pounds which when converted to kgs it becomes 14.969kgs.

Rounded to 15kgs.

Phenytoin loading dose is 15mg/kg

15mg×15 kg =225mgs. should be given

Prescribed is 0.5mg /kg/ min

225 mg ÷ 0.5 mg = 450 (0.5mgs present in 225mg)

450÷ 15 kg = 30 mgs/min

Pharmacy has 50mg Phenytoin/ml in a 5ml vial

225mgs÷ 50mg = 4.5 ml

So 4.5 ml should be loaded from vial

Normal saline 50ml + 4.5 ml drug = 54.5ml to be administered.

Per min 30 mgs should be given then, 225 mgs can be given in 225 ÷30 = 7.5 min

So , totally 225 mgs can be given in 8 min which is rounded to nearest value.